OpenStudy (anonymous):

Simplify: (4x − 6) + (3x + 6)

OpenStudy (jdoe0001):

just get rid of the parentheses and add away

OpenStudy (anonymous):

(4x-6) + (3x+6) = 4x - 6 + 3x + 6 = 7x +0 = 7x

OpenStudy (jdoe0001):

yeah